# Chain of Condition: Construct, Verify and Solve Conditions for Conditional Question Answering

### v1

@@ Abstract
Conditional question answering (CQA) is an important task that aims to find probable answers and identify conditions that need to be satisfied to support the answer. Existing approaches struggle with CQA due to two main challenges: (1) precisely identifying conditions and their logical relationship, and (2) verifying and solving the conditions. To address these challenges, we propose Chain of Condition, a novel prompting approach by firstly identifying all conditions and constructing their logical relationships explicitly according to the document, then verifying whether these conditions are satisfied, finally solving the logical expression by tools to indicate any missing conditions and generating the answer based on the resolved conditions. The experiments on two benchmark conditional question answering datasets shows chain of condition outperforms existing prompting baselines, establishing a new state-of-the-art. Furthermore, with backbone models like GPT-3.5-Turbo or GPT-4, it surpasses all supervised baselines with only few-shot settings. [^1] Codes will be released upon the acceptance of this paper.
@@
Jiuheng Lin1, Yuxuan Lai2,3, Yansong Feng1thanks: Corresponding author. 1 Wangxuan Institute of Computer Technology, Peking University, China 2 Department of Computer Science, The Open University of China 3 Engineering Research Center of Integration and Application of Digital Learning Technology, Ministry of Education linjiuheng@stu.pku.edu.cn    erutan@pku.org.cn    fengyansong@pku.edu.cn
@@ 1 Introduction
Conditional question answering (CQA) aims to answer questions where the information provided by the user may not be sufficient, therefore, additional conditions are necessary to imply the correctness of the answers  Saeidi et al. (2018); Min et al. (2020); Sun et al. (2021a); Dhingra et al. (2022). CQA is a challenging and promising task, which has been gaining increasing attention recently Sun et al. (2022); Du et al. (2023); Wang et al. (2023); Hussain et al. (2023); Puerto et al. (2024). An example is shown in Figure 1. The user asks for the amount of benefit she would receive, but according to the policy, the applicant must not claim for other benefits and has an unemployment certificate as prerequisite. These conditions are not mentioned in the user’s description, so a proper answer should include the amount of benefit as well as these unsatisfied conditions to ensure the eligibility.
@@
The major challenge for CQA is twofold. The first challenge is identifying conditions from the document and comprehending the relationships between them. In Figure 1, all conditions are highlighted, while the rest of the description concerning other aspects of Childbirth Benefit is irrelevant to the question. Moreover, there are different relationships between these conditions. For example, the conditions you are the birth parent and you didn’t claim other benefits must both be satisfied, whereas the conditions is the biological father of the child and has lived together for at least 3 years require at least one to be satisfied. Precisely identifying all conditions and understanding their relationships according to the document is a fundamental prerequisite for the CQA task, but existing approaches primarily build end-to-end systems that overlook this challenge Ainslie et al. (2020); Izacard and Grave (2021); Hussain et al. (2023). These methods take the whole document as input, train models to implicitly identify conditions and parse their relationships, and directly output the answer along with any missing conditions. Consequently, due to the limitations of implicit reasoning capabilities in models, these approaches struggle with questions involving multiple conditions and complex relationships. Besides, their solution path is impossible for users to interpret.
@@
Verifying and solving conditions with their logical relationships based on the user’s question is the second challenge Sun et al. (2022). Each condition may be satisfied, contradicted or ignored by the user, and conditions form a logical expression with each other based on their relationships. Solving this expression is necessary for accurately identifying all conditions missing from the user’s input. For example, in Figure 1, the conditions in green are satisfied by the user, while those in red are not mentioned. Moreover, although the condition has lived together for at least 3 years is not mentioned by the user, it has an "either" relationship with an already satisfied condition, is the biological father of the child. Therefore, the condition has lived together for at least 3 years is not necessary for user to satisfy. Determining the unnecessity of this condition requires correctly solving the logical expression of conditions. Previous works train models to verify and solve conditions simultaneously, requiring models to implicitly resolve the expression Du et al. (2023); Wang et al. (2023). This approach risks computational errors in solving expressions, and reduces the precision of predicting conditions.
@@
To address these challenges, we introduce Chain of Condition, a novel prompting framework for constructing, verifying, and solving conditions in the CQA task. Chain of condition include three main steps: first explicitly identifying all conditions from the document and constructing the logical expression of them according to the document, next verifying whether conditions have been satisfied by the user, finally solving the logical expression precisely by tools to indicate any missing conditions, and generating the appropriate answer based on the resolved conditions.
@@
We conduct experiments on two CQA benchmark datasets ConditionalQA Sun et al. (2021a) and ShARC Saeidi et al. (2018). The results show that chain of condition remarkably outperforms all prompting baselines. And with backbone models like GPT-3.5-Turbo or GPT-4, chain of condition even performs better than all supervised baselines.
@@
(1)We are the first to investigate prompting LLMs for conditional question answering task.
(2)We propose Chain of Condition, a construct, verify and solve prompting framework. Experiments show chain of condition outperforms existing prompting baselines on all benchmark datasets, establishing a new state-of-the-art. And with backbone models like GPT-3.5-Turbo or GPT-4, it even surpasses all fully supervised baselines with only few-shot settings.
(3)We decompose the CQA task for LLMs, creating a coherent and interpretable reasoning path that is easier for humans to understand.
@@ 2 Related Works
Prompting Methods for LLMs Large language models can be guided to solve tasks in a step-by-step manner Wei et al. (2023). For more complex reasoning tasks such as multi-hop QA Yang et al. (2018) or math problems Cobbe et al. (2021), previous works typically address them by decomposing the question into simple sub-questions for models to solve sequentially Yao et al. (2023); Press et al. (2023); Servantez et al. (2024); Tao et al. (2024). This decomposition reduces task difficulty and improves interpretability. Another approach to enhance performance on reasoning tasks is by combining LLMs with symbolic interpreters such as a Python runtime Chen et al. (2023); Gao et al. (2023); Lyu et al. (2023) or a SAT solver Ye et al. (2023). Compared to textual reasoning, program-based reasoning is executed accurately, thus achieving high-precision reasoning in complex questions. Furthermore, Puerto et al. (2024) compared the performance of these two approaches and showed that introducing code in the prompt may elicit the reasoning ability for the CQA task. In this work, we integrate the ideas of decomposing task and leveraging code interpreter into our chain of condition framework, enabling it to benefit from both approaches’ advantages. This not only improves interpretability but also increases precision.
@@ 3 Preliminary
We investigate the conditional question answering (CQA) task, where the answer is valid only when all missing conditions from the context are provided. Formally, the task’s input consists of the user’s question $Q$ and scenario $S$ paired with a reference document $D$. The answer should be inferred from the document. Unlike other QA tasks, the document in the CQA task contains numerous conditions $C=\{c_{1};c_{2};...;c_{n}\}\subset D$ that must be satisfied to obtain the answer. The complete output includes the answer along with any corresponding unmentioned conditions $A=(a,C^{\mathrm{(u)}})$, where $\{c_{1}^{\mathrm{(u)}};...;c_{k}^{\mathrm{(u)}}\}\subset C$ denotes the $i$-th unmentioned condition for answer $a$, and $k_{i}\geq 0$ denotes the total number of unmentioned conditions for the answer [^2] A few questions in ConditionalQA have multiple answers with conditions, and we leave the condition prediction for them as future work.. If there are no unmentioned conditions, then we categorize the answer as deterministic. Otherwise, we call it conditional, and all missing conditions should be listed simultaneously with the answer.
@@ 4 Methodology
We introduce Chain of Condition, a novel approach to guide Large Language Models (LLMs) for the conditional question answering task. Chain of condition pipeline is similar to the human process for dealing with such problems, and therefore demonstrates a more coherent and interpretable solving path for users to understand. The detailed process of this approach is illustrated in Figure 2.
@@
Chain of condition includes three steps: condition identification and logical construction, condition verification, expression solution and answer generation. These steps decompose the original CQA task into smaller sub-tasks, allowing the model to solve them sequentially. First, we identify all conditions and parse their logical relationships according to the document, forming a logical expression of conditions. Secondly, we verify each condition’s fulfillment in the user’s description. Finally, We take the verification results to solve the logical expression and identify all missing conditions by tool, and generate the answer with the aid of condition solutions. A prompt example for each step is in Appendix F.
4.1 Condition Identification and Logical Construction
The document $D$ contains a substantial amount of irrelevant information, so the first task is to filter this out and identify all relevant conditions $C=\{c_{1};c_{2};...;c_{n}\}\subset D$. By leveraging the powerful language comprehension capabilities of LLMs, we can instruct the model to directly identify the locations of conditions in the document.
@@ 4.1 Condition Identification and Logical Construction
In addition to identifying the conditions, it is crucial to arrange them in a particular logical order according to the document. We address this by instructing models to construct condition groups $G=\{G_{1};G_{2};...;G_{k}\}$, where $k$ is the total number of condition groups in the document. The $i$-th condition group $G_{i}$ is defined as $G_{i}(c_{1}^{i},c_{2}^{i},...,c_{n_{i}}^{i},o_{i})=(f(c_{1}^{i})\ o_{i}\ f(c_{2}^{i})\ o_{i}\ ...\ o_{i}\ f(c_{n_{i}}^{i}))$, where $\{c_{1}^{i},...,c_{n_{i}}^{i}\}$ are the conditions in the $i$-th group, $o_{i}\in\{\mathrm{and},\mathrm{or}\}$ is the logical operator connecting them and $f(c)$ represents the fulfillment of condition $c\in C$ that will be determined in the next step. We parse the model’s output to obtain $G$, ultimately forming a compositional logical expression of conditions $F(G_{1},...,G_{k},o)=((G_{1})\ o\ ...\ o\ (G_{k}))$, where $o\in\{\mathrm{and},\mathrm{or}\}$. The solving result of $G_{i}$ and $F$ are in $\{\bar{\textbf{d}},\bar{\textbf{c}}\}$, where $\bar{\textbf{d}}$ denote the deterministic answer and $\bar{\textbf{c}}$ conditional answer.
@@
Besides, the conditions in the document are usually free-form, making it difficult to thoroughly separate a condition from other irrelevant context. Simply truncating or slicing the document may result in incomplete conditions. For example, in Figure 2, the condition is the biological father of the child lacks a subject, which needs to be extracted from the previous sentence, your partner. This incompleteness could prevent the model from correctly understanding the meaning of the condition, consequently hindering its ability to accurately verify the fulfillment of the condition.
@@
To address this problem, we take context-based augmentation after identifying condition’s locations. This approach allows us to obtain a short paragraph for each condition, containing all the necessary additional information. Specifically, we employ two augmentation methods: leveraging the structural information of the document or using transcription. When the document has a certain structure, such as HTML tags for each paragraph, we use this to find the relevant context for augmentation. We take the entire subsection where the condition appears as the augmentation paragraph, ensuring it contains enough background information while being much shorter than the entire document. When there is no structural information available, we instruct the model to directly transcribe the condition based on the context.
@@ 4.2 Condition Verification
After acquiring all condition groups and augmenting each condition, we instruct the models to verify the fulfillment of each condition sequentially. This involves taking the question, scenario, and augmented condition as input, and leveraging the powerful reasoning capabilities of LLMs to determine the status of each condition. For each condition $c\in C$, the verification process can be formalized as determining the value of function $f(c)\in\{\tilde{\textbf{s}},\tilde{\textbf{c}},\tilde{\textbf{n}}\}$. Here, $\tilde{\textbf{s}}$ means the condition is satisfied by the user, $\tilde{\textbf{c}}$ means it is contradicted, and $\tilde{\textbf{n}}$ means the condition is not mentioned. Conditions that are either satisfied or contradicted lead to a deterministic answer, while conditions that are not mentioned result in a conditional answer. Therefore, the solving process of the expression $F$ can be seen as boolean operations on true/false values $f(c)$.
@@ 4.3 Expression Solution and Answer Generation
After obtaining the verification result $f(c)$ for each condition $c$, we need to recompose these results into groups $G$ and logical expression $F$, and follow the document to determine whether the answer is conditional or deterministic. And if the answer is conditional, all missing conditions should be listed along with the answer.
Traditionally, this is done by prompting models to implicitly reason and resolve the logical expression. However, recent studies have shown that even large language models struggle with logical or mathematical reasoning tasks Blair-Stanek et al. (2023). Therefore, a better solution is to offload the computation process to an external symbolic interpreter Chen et al. (2023); Gao et al. (2023); Lyu et al. (2023); Ye et al. (2023). In chain of condition, we use a Python interpreter to solve the logical expression $F(G_{1},...,G_{k},o)$. For conditional answers, we also identify all missing conditions $C^{\mathrm{(u)}}$ by $\{c_{i}^{j}|f(c_{i}^{j})=\tilde{\textbf{n}}\ ,\ G_{i}=\bar{\textbf{c}},\forall i\leq n_{i},\forall j\leq k\}$. This approach reduces model inference costs, improves precision, and enhances interpretability.
After obtaining the complete result for the conditions, we instruct the models to generate the answer. Since we have already verified each condition’s fulfillment, we can leverage this information for more accurate answer generation. Specifically, we add these conditions $c$ along with their fulfillment $f(c)$ into the prompt. This provides the model with straightforward information about the conditions, reducing the need to repeatedly infer their fulfillment from the document. Additionally, these conditions help the model locate relevant paragraphs about the question in the document.
@@
Table 3: Results of ConditionalQA compared with supervised baselines.

### v2

@@ Abstract
Conditional question answering (CQA) is an important task that aims to find probable answers and identify missing conditions. Existing approaches struggle with CQA due to two challenges: (1) precisely identifying necessary conditions and the logical relationship, and (2) verifying conditions to detect any that are missing. In this paper, we propose a novel prompting approach, Chain of condition, by first identifying all conditions and constructing their logical relationships explicitly according to the document, then verifying whether these conditions are satisfied, finally solving the logical expression to indicate any missing conditions and generating the answer accordingly. Experiments on two CQA benchmark datasets show our chain of condition outperforms existing prompting baselines, establishing a new state of the art. Furthermore, with only a few examples, our method can facilitate GPT-3.5-Turbo or GPT-4 to outperform all existing supervised models. [^1] Our code is open sourced at: https://github.com/Infinite-set/Chain-of-Condition
@@
Jiuheng Lin1, Yuxuan Lai2,3, Yansong Feng1thanks: Corresponding author. 1 Wangxuan Institute of Computer Technology, Peking University, China State Key Laboratory of General Artificial Intelligence, Peking University, China 2 Department of Computer Science, The Open University of China 3 Engineering Research Center of Integration and Application of Digital Learning Technology, Ministry of Education linjiuheng@stu.pku.edu.cn    erutan@pku.org.cn    fengyansong@pku.edu.cn
@@ 1 Introduction
Conditional question answering (CQA) aims to answer questions where the information provided by the user may not be sufficient, and any missing information should be requested from the user to determine the answer Saeidi et al. (2018); Min et al. (2020); Sun et al. (2021a); Dhingra et al. (2022); Ju et al. (2022); Zhang et al. (2023). CQA is a challenging and promising task, which has been gaining increasing attention recently Sun et al. (2022); Du et al. (2023); Wang et al. (2023); Hussain et al. (2023); Puerto et al. (2024). Figure 1 shows an example. The user asks for the amount of benefit she would receive, but according to the policy, the applicant must not claim for other benefits and has an unemployment certificate as prerequisite. These conditions are not mentioned in the user’s description, therefore it is necessary to remind the user to provide this missing information to determine the benefit amount, which is up to \$120000.
@@
The mains challenges for CQA are twofold. The first one is identifying all conditions from the document and analyzing their relationship. In Figure 1, all conditions are highlighted, while the majority of the description focuses on irrelevant aspects of Childbirth Benefit, making it difficult to accurately identify all relevant conditions. Moreover, there are specific relationships between conditions. For example, the conditions you are the birth parent and you didn’t claim other benefits must be satisfied simultaneously, whereas the conditions is the biological father of the child and has lived together for at least 3 years require at least one to be satisfied. Precisely identifying all conditions and their relationships according to the document is a prerequisite for the CQA task, but existing approaches primarily build end-to-end systems that overlook this challenge Ainslie et al. (2020); Izacard and Grave (2021); Hussain et al. (2023). These methods take the whole document as input, train models to implicitly identify conditions and parse their relationships, and directly output the answer along with any missing conditions. Consequently, due to limited reasoning abilities in existing models, these approaches struggle with questions involving multiple conditions and complex relationships among them. Besides, their solution path is implicit thus impossible for users to interpret.
@@
Verifying conditions and solving their logical relationships is the second challenge Sun et al. (2022). Each condition may be satisfied, contradicted or ignored by the user, and must be correctly verified. Furthermore, multiple conditions along with their logical relationship will form a logical expression, and solving this expression is necessary for accurately identifying all the conditions missing from the user’s input. For example, in Figure 1, the conditions in green are satisfied according to the query, while those in red are not mentioned. Moreover, although the condition has lived together for at least 3 years is not mentioned by the user, it has an "either" relationship with an already satisfied condition, is the biological father of the child. Therefore, the condition has lived together for at least 3 years is not necessary for user to satisfy. Determining the un-necessity of such condition requires correctly solving the logical expression of conditions. Previous works train models to verify and solve conditions simultaneously, requiring models to implicitly resolve the expression Du et al. (2023); Wang et al. (2023) and leading to computational errors and reduced precision.
@@
To address these challenges, we introduce Chain of Condition, a novel prompting framework for constructing, verifying and solving conditions in CQA tasks. Chain of condition include three main steps: first explicitly identifying all conditions and constructing the logical expression about the conditions according to the document, next verifying whether each condition has been satisfied by the user, finally solving the logical expression precisely through tool use to indicate missing conditions and generating the appropriate answer.
@@
We conduct experiments on two CQA benchmark datasets ConditionalQA Sun et al. (2021a) and ShARC Saeidi et al. (2018). The results show that chain of condition significantly outperforms all prompting baselines. And with backbone models like GPT-3.5-Turbo or GPT-4, our chain of condition even performs better than all supervised baselines.
@@
(1) We propose Chain of condition, an construct-verify-solve three-step prompting framework for conditional question answering. Experiments on benchmark datasets show our method outperforms existing prompting baselines, establishing a new state-of-the-art.
(2) Our proposed framework can enhance backbone models like GPT-3.5-Turbo or GPT-4 to surpasses all fully supervised baselines, with only a few examples.
(3) Our Chain of condition can create coherent and interpretable reasoning paths that are easier for humans to understand.
@@ 2 Related Works
Prompting Methods for LLMs Large language models can be guided to solve tasks in a step-by-step manner Wei et al. (2023). For more complex reasoning tasks such as multi-hop QA Yang et al. (2018) or math problems Cobbe et al. (2021), previous works typically address them by decomposing the question into simple sub-questions for models to solve sequentially Yao et al. (2023); Press et al. (2023); Yu et al. (2023); Servantez et al. (2024); Tao et al. (2024). This decomposition reduces task difficulty and improves interpretability. Another approach to enhance performance on reasoning tasks is by combining LLMs with symbolic interpreters such as a Python runtime Chen et al. (2023); Gao et al. (2023); Lyu et al. (2023) or a SAT solver Ye et al. (2023). Compared to textual reasoning, program-based reasoning is executed accurately, thus achieving high-precision reasoning in complex questions. Furthermore, Puerto et al. (2024) compared the performance of these two approaches and showed that introducing code in the prompt may elicit the reasoning ability for the CQA task. In this work, we integrate the ideas of decomposing task and leveraging code interpreter into our framework, enabling it to benefit from both approaches’ advantages. This not only improves interpretability but also increases precision.
@@ 3 Preliminary
We investigate the conditional question answering (CQA) task, where all missing condition should be requested from the user to determine the answer. Formally, the task’s input consists of the user’s question $Q$ and scenario $S$ providing some background information, and a reference document $D$ describing the policy being questioned. The answer should be inferred from the document. Unlike other QA tasks, the document in the CQA task contains many conditions =$\mathbb{C}=\{c_{1};c_{2};...;c_{n}\}\subset D$ that must be satisfied by the user. The complete output includes the answer along with any unmentioned conditions $A=(a,\mathbb{C}^{\mathrm{(u)}})$, where $\mathbb{C}^{\mathrm{(u)}}=\{c_{1}^{\mathrm{(u)}};...;c_{k}^{\mathrm{(u)}}\}\subset\mathbb{C}$ denotes the $i$-th unmentioned condition for answer $a$, and $k_{i}\geq 0$ denotes the total number of unmentioned conditions for it [^2] A few questions in ConditionalQA have multiple answers with conditions, and we leave the condition prediction for them as future work.. If there are no unmentioned conditions, then we categorize the answer as deterministic. Otherwise, we call it conditional, and all missing conditions should be listed simultaneously with the answer.
@@ 4 Methodology
We introduce Chain of Condition, a novel approach to guide Large Language Models (LLMs) for conditional question answering tasks. Our chain of condition pipeline is similar to the human process for dealing with such problems, thus can demonstrate a more coherent and interpretable solving path for users to read, as illustrated in Figure 2.
@@
Chain of condition includes three steps: condition identification and logical construction, condition verification, expression resolution and answer generation. Our framework decompose the original CQA task into smaller sub-tasks, allowing the model to solve them sequentially. First, we identify all conditions and recognize their logical relationships according to the document, forming a logical expression of conditions. Secondly, we verify each condition’s fulfillment in the user’s description. Finally, We take the verification results to solve the logical expression and identify all missing conditions through Python interpreter, and generate the answer by taking the condition solutions into account. A prompt example for each step is shown in Appendix G.
4.1 Condition Identification and Logical Relationship Construction
@@
The document $D$ contains a substantial amount of irrelevant information, so the first step is to filter this out and identify all relevant conditions =$\mathbb{C}=\{c_{1};c_{2};...;c_{n}\}\subset D$. In addition to identifying the conditions, it is crucial to arrange them in a particular structure according to the document. We address this by introducing condition groups $\mathbb{G}=\{G_{1};G_{2};...;G_{k}\}$, where $k$ represents the total number of condition groups in the document, and each group $G_{i}$ consists a set of nearby conditions sharing the same relationship. The $i$-th condition group $G_{i}$ is defined as $G_{i}(c_{1}^{i},c_{2}^{i},...,c_{n_{i}}^{i},o_{i})=(f(c_{1}^{i})\ o_{i}\ f(c_{2}^{i})\ o_{i}\ ...\ o_{i}\ f(c_{n_{i}}^{i}))$, where $\{c_{1}^{i},...,c_{n_{i}}^{i}\}$ are the conditions in the $i$-th group, $o_{i}\in\{\mathrm{AND},\mathrm{OR}\}$ is the logical operator connecting them and $f(c)$ represents the fulfillment of condition $c\in\mathbb{C}$ that will be verified in the next step. we instruct the model to directly identify the condition locations and their relationship in the document, then parse the model’s output to obtain $\mathbb{G}$, ultimately forming a compositional logical expression of conditions $F(G_{1},...,G_{k},o)=((G_{1})\ o\ ...\ o\ (G_{k}))$, where $o\in\{\mathrm{AND},\mathrm{OR}\}$. The solving result of $G_{i}$ and $F$ are in $\{\bar{\textbf{d}},\bar{\textbf{c}}\}$, where $\bar{\textbf{d}}$ denote the deterministic answer and $\bar{\textbf{c}}$ conditional answer. Detailed discussion of the solving process is in Section 4.3
@@
Besides, the conditions in the document are usually free-form, making it difficult to thoroughly separate a condition from other irrelevant context. Simply truncating the document may result in incomplete conditions. For example, in Figure 2, the condition is the biological father of the child lacks a subject your partner, which needs to be extracted from the previous sentence. This incompleteness could prevent the model from correctly understanding the meaning of the condition, consequently hindering its ability to accurately verify the fulfillment of the condition.
@@
To address this problem, we take context-based augmentation after identifying conditions’ locations. This approach allows us to obtain a short paragraph for each condition, containing all the necessary information. Specifically, when the document has a certain structure (e.g. subsections), we use this to find the relevant context for augmentation. We take the entire subsection where the condition appears as the augmentation paragraph, ensuring it contains enough background information while being much shorter than the entire document. When there is no structural information available, we instruct the model to directly summarize the condition based on the context.
@@ 4.2 Condition Verification
After acquiring all condition groups from $D$ and augmenting each condition, we instruct the models to sequentially verify the fulfillment of each condition based on $S$. This involves taking the question, scenario, and augmented condition as input, and leveraging LLMs to determine the fulfillment of conditions. For each condition $c\in\mathbb{C}$, the verification process can be formalized as determining the value of function $f(c)\in\{\tilde{\textbf{s}},\tilde{\textbf{c}},\tilde{\textbf{n}}\}$. Here, $\tilde{\textbf{s}}$ means the condition is satisfied by the user, $\tilde{\textbf{c}}$ means it is contradicted, and $\tilde{\textbf{n}}$ means the condition is not mentioned. Conditions that are either satisfied or contradicted lead to a deterministic answer and can be treated as true values, while unmentioned conditions result in a conditional answer and are treated as false. Therefore, solving the expression $F$ can be viewed as boolean operations on true/false values with special shortcut judgement. This process will be discussed further in Section 4.3.
@@ 4.3 Expression Solution and Answer Generation
After obtaining the verification result $f(c)$ for each condition $c$, we take these results into groups $\mathbb{G}$ and logical expression $F$ to determine whether the answer is conditional or deterministic. And if the answer is conditional, all missing conditions should be listed along with the answer.
Traditionally, this is done by prompting models to implicitly reason and resolve the logical expression. However, recent studies have shown that even large language models struggle with logical or mathematical reasoning tasks Blair-Stanek et al. (2023). Therefore, a better solution is to offload the computation process to an external symbolic interpreter Chen et al. (2023); Gao et al. (2023); Lyu et al. (2023); Ye et al. (2023).
In chain of condition, we use a Python interpreter to solve the logical expression $F(G_{1},...,G_{k},o)$. Specifically, we treat all "not mentioned" conditions $\{c|f(c)=\tilde{\textbf{n}}\}$ as false values, while "satisfied" or "contradicted" conditions $\{c|f(c)\in\{\tilde{\textbf{s}},\tilde{\textbf{c}}\}\}$ are considered true. Additionally, a "satisfied" condition linked by an "OR" relationship in its condition group or a "contradicted" condition with an "AND" relationship in its condition group $G_{i}$ would form a shortcut, allowing for direct verification of $G_{i}$. This process can be formally expressed as $\mathbb{G}^{\mathrm{sc}}=\left\{G_{i}\mid o_{i}=\mathrm{or},\,\exists c\in G_{i}\,\hskip 3.00003pt\text{s.t.}\,f(c)=\tilde{\textbf{s}}\right\}\cup\left\{G_{j}\mid o_{j}=\mathrm{and},\,\exists c\in G_{j}\,\hskip 3.00003pt\text{s.t.}\,f(c)=\tilde{\textbf{c}}\right\}$. Finally, a group $G_{i}$ is considered conditional (i.e., $G_{i}=\bar{\textbf{c}}$) if there exists unmentioned condition within it and no shortcut has been formed, formally $G_{i}\notin\mathbb{G}^{\mathrm{sc}},\ \exists c\in G_{i}\,\hskip 3.00003pt\text{s.t.}\,f(c)=\tilde{\textbf{c}}$, and otherwise, it is classified as deterministic. This reasoning process is automatically executed via Python using Boolean operations and special shortcut judgments.
Furthermore, we indicate all missing conditions, i.e., when a condition is not mentioned and all other conditions in the same group have not triggered any shortcut, formally $\mathbb{C}^{\mathrm{(u)}}=\{c_{i}^{j}|f(c_{i}^{j})=\tilde{\textbf{n}},G_{j}=\bar{\textbf{c}},G_{j}\notin\mathbb{G}^{\mathrm{sc}},\forall i\leq n_{j},\forall j\leq k\}$. If $\mathbb{C}^{\mathrm{(u)}}\neq\emptyset$, then the answer is conditional, and otherwise deterministic. By levaraging Python interpreter, we reduce model inference costs, improve precision, and enhance interpretability.
After obtaining the verification result for the conditions, we instruct the model to generate the answer. Since we have already verified each condition’s fulfillment, we can leverage this information for more accurate answer generation. Specifically, we add these conditions $c$ along with their fulfillment $f(c)$ into the prompt. This provides the model with straightforward information about the conditions, reducing the need to infer their fulfillment from the document.
@@
Table 3: Chain of condition vs. supervised baselines.